### v1

@@ 0.2.2 Robust Training
  • adversarial training [45, 85],
@@
Data augmentation involves the creation of synthetic examples through the application of diverse transformations or perturbations to the initial training data. These generated instances are then incorporated into the original dataset to enhance the training process. Adversarial training entails identifying worst-case examples at each epoch during the training phase and calculating an additional loss on these instances. State of the art adversarial training involve projected gradient descent algorithms such as FGSM [45] and PGD [85]. Certified training methods focus on providing mathematical guarantees about the model’s behaviour within certain bounds. Among them, we can name IBP training [46, 139] techniques, which impose intervals or bounds on the predictions or activations of the model, ensuring that the model’s output lies within a specific range with high confidence.
@@
Note that all techniques mentioned above can be categorised based on whether they primarily augment the data (such as data augmentation) or augment the loss function (as seen in adversarial, IBP and certified training). Augmenting the data tends to enhance generalisation and is efficient, albeit it may not help against the most severe adversarial attacks. Conversely, methods that manipulate the loss functions directly confront the toughest adversarial attacks but often come with higher computational costs. Ultimately, the choice between altering data or loss functions depends on the specific requirements of the application and the desired trade-offs between performance, computational complexity, and robustness guarantees.
NLP robustness.
@@
There exists a substantial body of research dedicated to enhancing the adversarial robustness of NLP systems [142, 121, 122, 75, 149, 150, 36]. These efforts aim to mitigate the vulnerability of NLP models to adversarial attacks and improve their resilience in real-world scenarios [121, 122] and mostly employ data augmentation techniques [40, 33]. In NLP, perturbations can occur at the character, word, or sentence level [24, 52, 20] and may involve deletion, insertion, swapping, flipping, substitution with synonyms, concatenation with characters or words, or insertion of numeric or alphanumeric characters [76, 39, 69]. For instance, in character level adversarial attacks, [9] introduces natural and synthetic noise to input data, while  [42, 72] identify crucial words within a sentence and perturbs them accordingly. For word level attacks, they can be categorised into gradient-based [76, 104], importance-based [51, 56], and replacement-based [3, 66, 98] strategies based on the perturbation method employed. In addition, in sentence level adversarial attacks, some attacks [53, 124] are created so that they do not impact the original label of the input and can be incorporated as a concatenation in the original text. In such scenarios, the expected behaviour from the model is to maintain the original output, and the attack can be deemed successful if the label/output of the model is altered. By augmenting the training data with these perturbed examples, models are exposed to a more diverse range of linguistic variations and potential adversarial inputs. This helps the models to generalise better and become more robust to different types of adversarial attacks. To help with this task, the NLP community has gathered a dataset of adversarial attacks named AdvGLUE [117], which aims to be a principled and comprehensive benchmark for NLP robustness measurements.
In this work we employ a PGD-based adversarial training as the method to increase the robustness and verifiability of our models, and we create character and word level perturbations as in [91] and sentence level perturbations with PolyJuice [132] and Vicuna [25].
0.2.3 Previous NLP Verification Approaches
@@
Method Verification algorithm Verification characteristics Datasets NLP perturbations Embeddings Architectures (# of parameters) Robust training
Ours
SMT-based, Abstract interpretation-based,
BaB-based
@@
RUARobot, Medical
@@
Sentence: S-BERT, S-GPT
FFNN ($10^{4}$)
PGD-based
Jia et al. (2019) [54]
IBP-based
Incomplete,
Imprecise,
@@
IMDB, SNLI
Word substitution
Word: GloVe
LSTM, CNN, BoW, Attention-based, ($10^{5}$)
IBP-based
Huang et al. (2019) [50]
IBP-based
Incomplete,
Imprecise,
@@
AGNews, SST
Char and word substitution
Word: GloVe
CNN ($10^{5}$)
IBP-based
Welbl et al. (2020) [126]
IBP-based
Incomplete,
Imprecise,
@@
SNLI, MNLI
Word deletion
Word: GloVe
Attention-based ($10^{5}$)
Data augmentation, random and beam search adversarial training, IBP-based
Zhang et al. (2021) [146]
IBP-based
Incomplete,
Imprecise,
@@
IMDB, SST, SST2
Word perturbations
Word: not specified
LSTM ($10^{5}$)
IBP-based
Wang et al. (2023) [123]
IBP-based
Incomplete,
Imprecise,
Deterministic
IMDB, YELP, SST2
Word substitution
Word: GloVe
CNN ($10^{5}$)
IBP-based: Embedding Interval Bound Constraint (EIBC) triplet loss
Ko et al. (2019) [61]
Abstract interpretation-based
Incomplete,
Imprecise,
Deterministic
CogComp QC
[…]
Word: not specified
RNN, LSTM ($10^{5}$)
Shi et al. (2020) [107]
Abstract interpretation-based
Incomplete,
Imprecise,
Deterministic
YELP, SST
[…]
Word: not specified
Transformer ($10^{6}$)
Du et al. (2021) [37]
Abstract interpretation-based
Incomplete,
Imprecise,
Deterministic
Rotten Tomatoes Movie Review, Toxic Comment
[…]
Word: GloVe
RNN, LSTM ($10^{5}$)
IBP-based
Bonaert et al. (2021) [14]
Abstract interpretation-based
Incomplete,
Imprecise,
Deterministic
SST, YELP
[…]
Word: not specified
Transformer ($10^{6}$)
Ye et al. (2020) [134]
Randomised smoothing ($\alpha=0.01,n=5000$)
Incomplete,
Imprecise,
@@
IMDB, Amazon
Word substitution
Word: GloVe
Transformer ($10^{8}$)
Data
@@
Wang et al. (2021) [120]
Differential privacy-based
Incomplete,
Imprecise,
@@
IMDB, AGNews
Word substitution
Word: GloVe
LSTM ($10^{5}$)
Data
augmentation
Zhao et al. (2022) [148]
Randomised smoothing ($\alpha=0.001,n=30050$)
Incomplete,
Imprecise,
Probabilistic
AGNews, SST
Word substitution
Word: GloVe
Transformer ($10^{8}$)
Data
@@
Zeng et al. (2023) [137]
Randomised smoothing ($\alpha=0.05,n=5000$)
Incomplete,
Imprecise,
@@
IMDB, YELP
Char and word substitution
Word: not specified
Transformer ($10^{8}$)
Data
@@
Ye et al. (2023) [135]
Randomised smoothing ($\alpha=0.001,n=9000$)
Incomplete,
Imprecise,
@@
IMDB, SST2, YELP, AGNews
Word substitution
Word: not specified
Transformer ($10^{8}$)
Data
@@
Zhang et al. (2023) [144]
Randomised smoothing ($\alpha=0.001,n=20000$)
Incomplete,
Imprecise,
@@
IMDB, Amazon, AGNews
Word perturbations
Word: GloVe
LSTM, Transformer ($10^{8}$)
Data
@@
Zhang et al. (2023) [147]
Randomised smoothing ($\alpha=0.05,n=5000$)
Incomplete,
Imprecise,
@@
SST2, AGNews
Word perturbations
Word: not specified
Transformer ($10^{9}$)
@@
Table 1: Summary of the main features of the existing NLP verification approaches (first half) and randomised smoothing approaches (second half). In bold are SoA methods.
@@
Although DNN verification studies have predominantly focused on computer vision, there is a growing body of research exploring the verification of NLP. This research can be categorised into three main approaches: IBP, abstract interpretation, and randomised smoothing. Table 1 shows a comparison of these approaches. To the best of our knowledge, this paper is the first one to use an SMT-based verifier for this purpose, and compare it with an abstract interpretation-based verifier on the same benchmarks.
Verification via Interval Bound Propagation.
@@
The first technique successfully adapted from the computer vision domain for verifying NLP models was the IBP. In the NLP approaches, IBP is used for both training and verification. Its aim is to minimise the upper bound on the maximum difference between the classification boundary and the input perturbation region by augmenting the loss function. This facilitates the minimisation of the perturbation region in the last layer, ensuring it remains on one side of the classification boundary. As a result, the adversarial region becomes tighter and can be considered certified robust. Notably, Jia et al. [54] proposed certified robust models on word substitutions in text classification. The authors employed IBP to optimise the upper bound over perturbations, providing an upper bound over the discrete set of perturbations in the word vector space. Furthermore, Huang et al. [50] introduced a verification and verifiable training method for neural networks in NLP, proposing a tighter over-approximation in the form of a ‘simplex’ in the embedding space for input perturbations. To make the network verifiable, they defined the convex hull of all the original unperturbed inputs as a space of perturbations. By employing the IBP algorithm, they generated robustness bounds for each neural network layer. Later on, Welbl et al. [126] differentiated from the previous approaches by using IBP to address the under-sensitivity issue. They designed and formally verified the ‘under-sensitivity specification’ that a model should not become more confident as arbitrary subsets of input words are deleted. Recently, Zhang et al. [146] introduced Abstract Recursive Certification (ARC) to verify the robustness of LSTMs. ARC defines a set of programmatically perturbed string transformations to construct a perturbation space. By memorising the hidden states of strings in the perturbation space that share a common prefix, ARC can efficiently calculate an upper bound while avoiding redundant hidden state computations. Finally, Wang et al. [123] improved on the work of Jia et al. by introducing Embedding Interval Bound Constraint (EIBC). EIBC is a new loss that constraints the word embeddings in order to tighten the IBP bounds.
@@
Another popular verification technique applied to various NLP models is based on abstract interpretation. Abstract interpretation was first developed by Cousot and Cousot [29] in 1977. It formalises the idea of abstraction of mathematical structures, in particular those involved in the specification of properties and proof methods of computer systems [28] and it has since been used in many applications [30]. Specifically, for DNN verification, this technique can model the behaviour of a network using an abstract domain that captures the possible range of values the network can output for a given input. This abstract domain can then be used to reason about the network’s behaviour under different conditions, such as when the network receives inputs that are adversarially perturbed. One notable contribution in this area is POPQORN [61], which is the first work that gives robustness guarantees for RNN-based networks. They handle the challenging non-linear activation functions of complicated RNN structures (like LSTMs and GRUs) by bounding them with linear functions. Later on, Du et al. improve on POPQORN by introducing Cert-RNN [37], a robust certification framework for RNNs that overcomes the limitations of POPQORN. The framework maintains inter-variable correlation and accelerates the non-linearities of RNNs for practical uses. Cert-RNN utilised Zonotopes [44] to encapsulate input perturbations and can verify the properties of the output Zonotopes to determine certifiable robustness. This results in improved precision and tighter bounds, leading to a significant speedup compared to POPQORN. Differently, Shi et al. [107] focus on transformers with self-attention layers. They developed a verification algorithm that can provide a lower bound to ensure the probability of the correct label is consistently higher than that of the incorrect labels. Analogously, Bonaert et al. [14] propose DeepT, a certification method for large transformers. It is specifically designed to verify the robustness of transformers against synonym replacement-based attacks. DeepT employes multi-norm Zonotopes to achieve larger robustness radii in the certification and can work with networks much larger than Shi et al.
[…]
Verification via Randomised Smoothing.
@@
Randomised smoothing [27] is another technique for verifying the robustness of deep language models that has recently grown in popularity due to its scalability [134, 120, 148, 137, 135, 144, 147]. Its basic idea is to leverage randomness during inference to create a smoothed classifier that is more robust to small perturbations in the input. This technique can also be used to give certified guarantees against adversarial perturbations within a certain radius. Generally, randomized smoothing begins by training a regular neural network on a given dataset. During the inference phase, to classify a new sample, noise is randomly sampled from the predetermined distribution multiple times. These instances of noise are then injected into the input, resulting in noisy samples. Subsequently, the base classifier generates predictions for each of these noisy samples. The final prediction is determined by the class with the highest frequency of predictions, thereby shaping the smoothed classifier. To certify the robustness of the smoothed classifier against adversarial perturbations within a specific radius centered around the input, randomised smoothing calculates the likelihood of agreement between the base classifier and the smoothed classifier when noise is introduced to the input. If this likelihood exceeds a certain threshold, it indicates the certified robustness of the smoothed classifier within the radius around the input.

### v3

@@ 2.3 Robust Training
  • adversarial training [71, 20] including property-driven training [74, 75],
@@
Data augmentation involves the creation of synthetic examples through the application of diverse transformations or perturbations to the initial training data. These generated instances are then incorporated into the original dataset to enhance the training process. Adversarial training entails identifying worst-case examples at each epoch during the training phase and calculating an additional loss on these instances. State of the art adversarial training involves projected gradient descent algorithms such as FGSM [71] and PGD [20]. Certified training methods focus on providing mathematical guarantees about the model’s behaviour within certain bounds. Among them, we can name IBP training [48, 76] techniques, which impose intervals or bounds on the predictions or activations of the model, ensuring that the model’s output lies within a specific range with high confidence.
@@
Note that all techniques mentioned above can be categorised based on whether they primarily augment the data (such as data augmentation) or augment the loss function (as seen in adversarial, IBP and certified training). Augmenting the data tends to be efficient, although it may not help against stronger adversarial attacks. Conversely, methods that manipulate the loss functions directly are more resistant to strong adversarial attacks but often come with higher computational costs. Ultimately, the choice between altering data or loss functions depends on the specific requirements of the application and the desired trade-offs between performance, computational complexity, and robustness guarantees.
2.4 NLP robustness
There exists a substantial body of research dedicated to enhancing the adversarial robustness of NLP systems [79, 80, 81, 82, 83, 84, 85]. These efforts aim to mitigate the vulnerability of NLP models to adversarial attacks and improve their resilience in real-world scenarios [80, 81] and mostly employ data augmentation techniques [86, 87]. In NLP, we can distinguish perturbations based on three main criteria:
  • where and how the perturbations occur,
  • whether they are altered automatically using some defined rules (vs. generated by humans or LLMs) and
  • whether they are adversarial (as opposed to random).
@@ 2.4 NLP robustness
In particular, perturbations can occur at the character, word, or sentence level [88, 89, 90] and may involve deletion, insertion, swapping, flipping, substitution with synonyms, concatenation with characters or words, or insertion of numeric or alphanumeric characters [91, 92, 93]. For instance, in character level adversarial attacks, Belinkov et al. [94] introduce natural and synthetic noise to input data, while Gao et al. [95] and Li et al. [96] identify crucial words within a sentence and perturb them accordingly. For word level attacks, they can be categorised into gradient-based [91, 97], importance-based [98, 99], and replacement-based [100, 101, 102] strategies, based on the perturbation method employed. Moreover, Moradi et al. [103] introduce rule-based non-adversarial perturbations at both the character and word levels. Their method simulates various types of noise typically caused by spelling mistakes, typos, and other similar errors. In sentence level adversarial attacks, some perturbations [104, 105] are created so that they do not impact the original label of the input and can be incorporated as a concatenation in the original text. In such scenarios, the expected behaviour from the model is to maintain the original output, and the attack can be deemed successful if the label/output of the model is altered. Additionally, non-rule-based sentence perturbations can be obtained through prompting LLMs [21, 14] to generate rephrasing of the inputs. By augmenting the training data with these perturbed examples, models are exposed to a more diverse range of linguistic variations and potential adversarial inputs. This helps the models to generalise better and become more robust to different types of adversarial attacks. To help with this task, the NLP community has gathered a dataset of adversarial attacks named AdvGLUE [106], which aims to be a principled and comprehensive benchmark for NLP robustness measurements.
In this work we employ a PGD-based adversarial training as the method to enhance the robustness and verifiability of our models against gradient-based adversarial attacks. For non-adversarial perturbations, we create rule-based perturbations at the character and word level as in Moradi et al. [103] and non-rule-based perturbations at the sentence level using PolyJuice [21] and Vicuna [14]. We thus cover most combinations of the three choices above (bypassing only human-generated adversarial attacks as there is no sufficient data to admit systematic evaluation which is important for this study).
@@
Method Datasets NLP perturbations Embeddings Architectures (# of parameters) Robust training Verification algorithm Verification characteristics
Ours RUARobot,
Medical
[…]
@@
Jia et al. (2019) [17] IMDB, SNLI Word substitution Word: GloVe LSTM, CNN, BoW, Attention-based, ($10^{5}$) IBP-based loss function augmentation Abstract Interpretation IBP Incomplete,
Deterministic
Huang et al. (2019) [18] AGNews, SST Char and word substitution Word: GloVe CNN ($10^{5}$) IBP-based loss function augmentation Abstract Interpretation IBP Incomplete,
Deterministic
Welbl et al. (2020) [107] SNLI, MNLI Word deletion Word: GloVe Attention-based ($10^{5}$) Data augmentation, random and beam search adversarial training, IBP-based Abstract Interpretation IBP Incomplete,
Deterministic
Zhang et al. (2021) [19] IMDB, SST, SST2 Word perturbations Word: not specified LSTM ($10^{5}$) IBP-based loss function augmentation Abstract Interpretation IBP Incomplete,
Deterministic
Wang et al. (2023) [108] IMDB, YELP, SST2 Word substitution Word: GloVe CNN ($10^{5}$) IBP-based: Embedding Interval Bound Constraint (EIBC) triplet loss Abstract Interpretation IBP Incomplete,
Deterministic
@@
[…]
@@
[…]
@@
[…]
@@
Table 1: Summary of the main features of the existing NLP verification approaches. In bold are state-of-the-art methods.
Method Datasets NLP perturbations Embeddings Architectures (# of parameters) Robust training Verification algorithm Verification characteristics
Ye et al. (2020) [112] IMDB, Amazon Word substitution Word: GloVe Transformer ($10^{8}$) Data
augmentation
Randomised smoothing ($\alpha=0.01,n=5000$) Incomplete,
@@
Wang et al. (2021) [113] IMDB, AGNews Word substitution Word: GloVe LSTM ($10^{5}$) Data
@@
Differential privacy Incomplete,
@@
Zhao et al. (2022) [114] AGNews, SST Word substitution Word: GloVe Transformer ($10^{8}$) Data
@@
Randomised smoothing ($\alpha=0.001,n=30050$) Incomplete,
@@
Zeng et al. (2023) [115] IMDB, YELP Char and word substitution Word: not specified Transformer ($10^{8}$) Data
@@
Randomised smoothing ($\alpha=0.05,n=5000$) Incomplete,
@@
Ye et al. (2023) [116] IMDB, SST2, YELP, AGNews Word substitution Word: not specified Transformer ($10^{8}$) Data
@@
Randomised smoothing ($\alpha=0.001,n=9000$) Incomplete,
@@
Zhang et al. (2023) [117] IMDB, Amazon, AGNews Word perturbations Word: GloVe LSTM, Transformer ($10^{8}$) Data
@@
Randomised smoothing ($\alpha=0.001,n=20000$) Incomplete,
@@
Zhang et al. (2023) [118] SST2, AGNews Word perturbations Word: not specified Transformer ($10^{9}$) - Randomised smoothing ($\alpha=0.05,n=5000$) Incomplete,
Probabilistic
@@
Table 2: Summary of the main features of the existing randomised smoothing approaches.
2.5 Datasets and Use Cases Used in NLP Verification
Existing NLP verification datasets. Table 3 summarises the main features and tasks of the datasets used in NLP verification. Despite their diverse origins and applications, the datasets in the literature are usually binary or multi-class text classification problems. Furthermore, datasets can be sensitive to perturbations, i.e. perturbations can have non-trivial impact on label consistency. For example, Jia et al. [17] use IBP with the SNLI [119] [^3] A semantic inference dataset that labels whether one sentence entails, contradicts or is neutral to another sentence. dataset (see Tables 1 and 3) to show that word perturbations (e.g. ‘good’ to ‘best’) can change whether one sentence entails another. Some works such as Jia et al. [17] try to address this label consistency, while others do not.
Additionally, we find that the previous research on NLP verification does not utilise safety critical datasets (which strongly motivates the choice of datasets in alternative verification domains), with the exception of Du et al. [110] that use the Toxic Comment dataset [120]. Other papers do not provide detailed motivation as to why the dataset choices were made, however it could be due to the datasets being commonly used in NLP benchmarks (IMDB etc.).
Dataset Safety Critical Category Tasks Size Classes
IMDB [121] $\times$ Sentiment analysis Document-level and sentence-level classification 25,000 2
SST [122] $\times$ Sentiment analysis Sentiment classification, hierarchical sentiment classification, sentiment span detection 70,042 5
SST2 [122] $\times$ Sentiment analysis Sentiment classification 70,042 2
YELP [123] $\times$ Sentiment analysis Sentiment classification 570,771 2
Rotten Tomatoes Movie Review [124] $\times$ Sentiment analysis Sentiment classification 48,869 3/4
Amazon [125] $\times$ Sentiment analysis Sentiment classification, aspect-based sentiment analysis 34,686,770 5
SNLI [119] $\times$ Semantic inference Natural language inference, semantic similarity 570,152 3
MNLI [126] $\times$ Semantic inference Natural language inference, semantic similarity, generalisation 432,702 3
AGNews [127] $\times$ Text analysis Text classification, sentiment classification 127,600 4
CogComp QC [128] $\times$ Text analysis Question classification, semantic understanding 15,000 6/50
Toxic Comment [120] $\sqrt{}$ Text analysis Toxic comment classification, fine-grained toxicity analysis, bias analysis 18,560 6
2.5.1 Datasets Proposed in This Paper
In this paper, we focus on two existing datasets that model safety-critical scenarios. These two datasets have not previously been applied or explored in the context of NLP verification. Both are driven by real-world use cases of safety-critical NLP applications, i.e. applications for which law enforcement and safety demand formal guarantees of “good” DNN behaviour.
Chatbot Disclosure (R-U-A-Robot Dataset [129]). The first case study is motivated by new legislation which states that a chatbot must not mislead people about its artificial identity [11, 10]. Given that the regulatory landscape surrounding NLP models (particularly LLMs and generative AI) is rapidly evolving, similar legislation could be widespread in the future – with recent calls for the US Congress to formalise such disclosure requirements [130]. The prohibition on deceptive conduct act may apply to the outputs generated by NLP systems if used commercially [131], and at minimum a system must guarantee a truthful response when asked about its agency [129, 132]. Furthermore, the burden of this should be placed on the designers of NLP systems, and not on the consumers.
Our first safety critical case is the R-U-A-Robot dataset [129], a written English dataset consisting of 6800 variations on queries relating to the intent of ‘Are you a robot?’, such as ‘I’m a man, what about you?’. The dataset was created via a context-free grammar template, crowd-sourcing and pre-existing data sources. It consists of 2,720 positive examples (where given the query, it is appropriate for the system to state its non-human identity), 3,400 negative examples and 680 ‘ambiguous-if-clarify’ examples (where it is unclear whether the system is required to state its identity). The dataset was created to promote transparency which may be required when the user receives unsolicited phone calls from artificial systems. Given systems like Google Duplex [133], and the criticism it received for human-sounding outputs [134], it is also highly plausible for the user to be deceived regarding the outputs generated by other NLP-based systems [131]. Thus we choose this dataset to understand how to enforce such disclosure requirements. We collapse the positive and ambiguous examples into one label, following the principle of ‘better be safe than sorry’, i.e. prioritising a high recall system.
Medical Safety Dataset. Another scenario one might consider is that inappropriate outputs of NLP systems have the potential to cause harm to human users [13]. For example, a system may give a user false impressions of its ‘expertise’ and generate harmful advice in response to medically related user queries [7]. In practice it may be desirable for the system to avoid answering such queries. Thus we choose the Medical safety dataset [12], a dataset consisting of 2,917 risk-graded medical and non-medical queries (1,417 and 1,500 examples respectively). The dataset was constructed via collecting questions posted on reddit, such as r/AskDocs. The medical queries have been labelled by experts and crowd annotators for both relevance and levels of risk (i.e. non-serious, serious to critical) following established World Economic Forum (WEF) risk levels designated for chatbots in healthcare [135]. We merge the medical queries of different risk-levels into one class, given the high scarcity of the latter two labels to create an in-domain/out-of-domain classification task for medical queries. Additionally, we consider only the medical queries that were labelled as such by expert medical practitioners. Thus this dataset will facilitate discussion on how to guarantee a system recognises medical queries, in order to avoid generating medical output.
2.6 Previous NLP Verification Approaches
@@
Although DNN verification studies have predominantly focused on computer vision, there is a growing body of research exploring the verification of NLP. This research can be categorised into three main approaches: using IBP, zonotopes, and randomised smoothing. Tables 1 and 2 show a comparison of these approaches. To the best of our knowledge, this paper is the first one to use an SMT-based verifier for this purpose, and compare it with an abstract interpretation-based verifier on the same benchmarks.
@@
NLP Verification via Interval Bound Propagation. The first technique successfully adopted from the computer vision domain for verifying NLP models was the IBP. IBP was used for both training and verification with the aim to minimise the upper bound on the maximum difference between the classification boundary and the input perturbation region. It was achieved by augmenting the loss function with a term that penalises large perturbations. Specifically, IBP incorporates interval bounds during the forward propagation phase, adding a regularisation term to the loss function that minimises the distance between the perturbed and unperturbed outputs. This facilitated the minimisation of the perturbation region in the last layer, ensuring it remained on one side of the classification boundary. As a result, the adversarial region becomes tighter and can be considered certifiably robust. Notably, Jia et al. [17] proposed certified robust models on word substitutions in text classification. The authors employed IBP to optimise the upper bound over perturbations, providing an upper bound over the discrete set of perturbations in the word vector space. Similarly, POPQORN[109] introduced robustness guarantees for RNN-based networks by handling the non-linear activation functions of complicated RNN structures (like LSTMs and GRUs) using linear bounds. Later, Shi et al.[15] developed a verification algorithm for transformers with self-attention layers. This algorithm provides a lower bound to ensure the probability of the correct label remains consistently higher than that of the incorrect labels. Furthermore, Huang et al. [18] introduced a verification and verifiable training method with a tighter over-approximation in style of the Simplex algorithm [28]. To make the network verifiable, they defined the convex hull of all the original unperturbed inputs as a space of perturbations. By employing the IBP algorithm, they generated robustness bounds for each neural network layer. Later on, Welbl et al. [107] differentiated from the previous approaches by using IBP to address the under-sensitivity issue. They designed and formally verified the ‘under-sensitivity specification’ that a model should not become more confident as arbitrary subsets of input words are deleted. Recently, Zhang et al. [19] introduced Abstract Recursive Certification (ARC) to verify the robustness of LSTMs. ARC defines a set of programmatically perturbed string transformations to construct a perturbation space. By memorising the hidden states of strings in the perturbation space that share a common prefix, ARC can efficiently calculate an upper bound while avoiding redundant hidden state computations. Finally, Wang et al. [108] improved on the work of Jia et al. by introducing Embedding Interval Bound Constraint (EIBC). EIBC is a new loss that constraints the word embeddings in order to tighten the IBP bounds.
@@
NLP Verification via Propagating Zonotopes. Another popular verification technique applied to various NLP models is based on propagating zonotopes, which produces tighter bounds then IBP methods. One notable contribution in this area is Cert-RNN [110], a robust certification framework for RNNs that overcomes the limitations of POPQORN. The framework maintains inter-variable correlation and accelerates the non-linearities of RNNs for practical uses. Cert-RNN utilised zonotopes [136] to encapsulate input perturbations and can verify the properties of the output zonotopes to determine certifiable robustness. This results in improved precision and tighter bounds, leading to a significant speedup compared to POPQORN. Analogously, Bonaert et al. [111] propose DeepT, a certification method for large transformers. It is specifically designed to verify the robustness of transformers against synonym replacement-based attacks. DeepT employs multi-norm zonotopes to achieve larger robustness radii in the certification and can work with networks much larger than Shi et al.
Methods that propagate zonotopes produce much tighter bounds than IBP-based methods, which can be used with deeper networks. However, they use geometric methods and do not take into account semantic considerations (e.g. do not use semantic perturbations).
@@
NLP Verification via Randomised Smoothing. Randomised smoothing [137] is another technique for verifying the robustness of deep language models that has recently grown in popularity due to its scalability [112, 113, 114, 115, 116, 117, 118]. The idea is to leverage randomness during inference to create a smoothed classifier that is more robust to small perturbations in the input. This technique can also be used to give certified guarantees against adversarial perturbations within a certain radius. Generally, randomized smoothing begins by training a regular neural network on a given dataset. During the inference phase, to classify a new sample, noise is randomly sampled from the predetermined distribution multiple times. These instances of noise are then injected into the input, resulting in noisy samples. Subsequently, the base classifier generates predictions for each of these noisy samples. The final prediction is determined by the class with the highest frequency of predictions, thereby shaping the smoothed classifier. To certify the robustness of the smoothed classifier against adversarial perturbations within a specific radius centered around the input, randomised smoothing calculates the likelihood of agreement between the base classifier and the smoothed classifier when noise is introduced to the input. If this likelihood exceeds a certain threshold, it indicates the certified robustness of the smoothed classifier within the radius around the input.